# Hydrostatic mass bias for galaxy groups and clusters in the FLAMINGO simulations

### v1

@@ 1 Introduction
Alternatively, the mass can be estimated using the hydrostatic mass inferred from X-ray observations. Assuming an emission model, the spatially resolved X-ray spectrum can be converted into density and temperature profiles (e.g. Arnaud, 1996). Under the assumption of hydrostatic equilibrium, fits to these profiles with an analytic function can be used to compute the enclosed mass (see e.g. Sarazin, 1988).
In practice, hydrostatic mass measurements pose a significant challenge. To obtain the mass enclosed within a sphere, the density and temperature profiles have to be three-dimensional, which means observations have to be deprojected from the plane of the sky in which they are observed (see e.g. Turner et al., 2024, for a comprehensive description of the process). Unlike in the case for weak lensing mass estimates, the cluster shape only has a 1 per cent effect on the mass estimate (Buote & Humphrey, 2012). However, temperature inhomogeneities within a single radial bin of the profile can have a 10 per cent effect on the mass, due to clumpy cold gas biasing the mean temperature low (Rasia et al., 2014; Henson et al., 2017).
@@
The hydrostatic masses inferred from X-ray observations are thus found to be biased. One method of determining the hydrostatic mass bias is by comparing the X-ray inferred masses ($M_{\rm 500c,X}$) with weak lensing masses ($M_{\rm 500c,WL}$). Using this approach, the XXL cluster survey found a value of $M_{\rm X}/M_{\rm WL}=0.72\pm 0.08$ for $2\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ (Eckert et al., 2016), and the Canadian cluster project obtained $M_{\rm X}/M_{\rm WL}=0.76\pm 0.06$ for $4\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<2\times 10^{15}~{}\mathrm{M_{\odot}}$ (Hoekstra et al., 2015). However, much weaker biases of $M_{\rm X}/M_{\rm WL}=0.95\pm 0.05$ for $2\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ were found by Smith et al. (2016) and later $M_{\rm X}/M_{\rm WL}=0.84\pm 0.095$ for $3\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ by Wicker et al. (2023).
@@ 2.1 The FLAMINGO simulations
FLAMINGO (Full-hydro Large-scale structure simulations with All-sky Mapping for the Interpretation of Next Generation Observations) is a large suite of hydrodynamical cosmological simulations, covering large cosmic volumes. The flagship run comprises a region of $\mathrm{(2.8~{}Gpc)^{3}}$, which is sufficiently large for statistical studies of galaxy groups and clusters. The simulations are described in detail in Schaye et al. (2023). A unique feature is the machine learning-aided calibration of the subgrid prescriptions for stellar and AGN feedback to the observed $z=0$ galaxy stellar mass function and $z=0.1-0.3$ cluster gas fractions (Kugel et al., 2023). Variations on the fiducial model are made by shifting the observed gas fractions up and down by a multiple of their uncertainty ($\sigma$), and recalibrating the model to fit those new data points. FLAMINGO also includes models that vary the galaxy mass function or the cosmology. We will not use the latter here as we found that the cosmology variations do not affect the bias.
@@ 2.4 Fitting the thermodynamic profiles
where $\alpha$, $\beta$, $\gamma$, $\epsilon$ are free parameters, as well as the normalisation of the density profile $\rho(0)$, the core radius $r_{\rm c}$, and the scale radius $r_{\rm s}$. The scale radius ($r_{\rm s}$) allows for a second characteristic transition radius in the profile. Though this model has 7 free parameters, $\gamma$ is conventionally set to $\gamma=3$, which we adopt throughout this work.
@@
an eight-parameter function, where $T_{\rm 0}$ represents the temperature normalisation, $T_{\rm min}$ the minimum temperature of the core, $r_{\rm c}$ the core radius, $r_{\rm t}$ the transition radius, which, along with $\alpha$, $\alpha_{\rm c}$, $\beta$, and $\gamma$, are free parameters.
@@ 3.1 Dependence on mass
Observational determinations of the bias, shown as the red data points, scatter widely below $10^{15}~{}\mathrm{M_{\odot}}$. We extract the data points from the observational literature by taking the highest and lowest mass in each observed sample and applying the power-laws, with errors, fitted by the respective authors at those masses. These data points are obtained using either $\beta$- or double-$\beta$-profiles to fit the density profile, and the Vikhlinin (2006) prescription for the temperature profile. Our results agree very well with Wicker et al. (2023) at $3\times 10^{14}~{}\mathrm{M_{\odot}}$, and have a very similar mass trend. In the study from Eckert et al. (2016) the inverse mass trend is found, where the bias decreases with increasing mass, though the magnitude is very similar to our result and the trend is weak. Hoekstra et al. (2015) find a much stronger mass dependence, because of which they find a weaker bias at their low-mass end, but a stronger bias at the highest masses. We conclude that the FLAMINGO hydrostatic biases are in good agreement with the observations, but that the observational constraints are rather uncertain for all but the highest masses.
@@ 3.3 X-ray versus volume weighted averaging
Because hydrostatic masses are inferred from spherically averaged profiles, any deviations from spherical symmetry will complicate their interpretation. The presence of structure will lead to deviations between thermodynamic profiles computed using different weighting schemes. Our approach of using X-ray weighted thermodynamic profiles in the hydrostatic mass inference is appropriate when comparing with results from X-ray observations, but to measure the true dynamical state of the cluster volume weighting is the natural choice. If the gas is clumpy, then X-ray weighting will favour denser gas, provided it is not sufficiently dense to cool to low temperatures, because the emissivity scales with the square of the density.
In Fig. 2, we compare the hydrostatic mass bias measured from X-ray- and volume weighted profiles. It is clear that the X-ray-weighted profiles result in a significantly stronger, and more mass dependent bias. As a large mass bias is driven by shallow gradients, this indicates that the thermal pressure gradient in X-ray-weighted gas is shallower. We conclude that for $M_{\text{500c}}\gg 10^{13}\,\text{M}_{\odot}$ most of the hydrostatic mass bias is due to X-rays being a biased tracer of the gas rather than due to true deviations from hydrostatic equilibrium.
@@ 3.4 Model variations
We note that even the most extreme feedback variations cannot reach the mass-dependent bias value obtained by Hoekstra et al. (2015) for the low-mass end of their sample. In general, the difference between the models is of the same order as the error on the observationally determined bias.
@@ 4.1 Fit quality
Figure 7: The logarithmic derivatives of the luminosity-weighted modified-$\beta$ density profiles and the luminosity-weighted temperature profiles for the objects with the 5 per cent lowest hydrostatic masses (black) or 5 per cent highest masses (red) in 5 narrow true mass bins. The legend in each panel indicates the median bias values of each subset. The two sets of contours are consistently offset and can be connected by moving almost orthogonal to the thin black diagonal lines which correspond to a constant pressure derivative, showing that the difference in bias could be explained by differences in the profile derivatives at a constant temperature (see Eq. 2).
@@ 4.2 Drivers of outlying hydrostatic mass
Fig. 7 reveals a separation between the two groups in the $\mathrm{dlog}\rho$-$\mathrm{dlog}T$ plane. Moving along the thin diagonal lines, the bias would change due to a different temperature at a fixed sum of derivatives (i.e. a fixed pressure derivative, see Eq. 2), whereas moving perpendicular to those lines the bias changes at fixed temperature, but due to changing derivatives.
@@
From Fig. 7 it can be seen that the two subsets can be connected by moving almost perpendicularly to the thin lines, indicating that the change in inferred hydrostatic mass could be explained by a change in the pressure gradient at fixed temperature. This agrees with results from Ansarifard et al. (2020), who provide a partial empircal correction to the hydrostatic mass bias by using the gradient of the density profile. To confirm this, we examine the temperatures at $r_{\rm 500c}$ for these two groups of outliers. Fig. 8 illustrates that objects exhibiting low and high inferred hydrostatic masses have comparable temperatures within identical narrow true mass bins. Consequently, we infer that the scatter in the bias at a fixed mass and at fixed goodness-of-fit can be accounted for by differences in the derivatives of the temperature and density profiles, as depicted in Fig. 7, indicating steeper (shallower) pressure profiles for the largest (smallest) hydrostatic masses at fixed true mass.
@@ 4.4 Energy balance
Martizzi & Agrusa (2016), who simulated clusters in the $M_{\rm 500c}\approx 1-3\times 10^{14}~{}\mathrm{M_{\odot}}$ range, found that, using volume weightig, at $r_{\rm 500c}$ non-thermal energy can fully account for the hydrostatic bias. We find that in this mass range, the non-thermal energy in FLAMINGO is indeed sufficient to achieve virial equilibrium if we ignore external pressure. Our estimate of $\approx 10$ per cent non-thermal energy is furthermore in line with what is found in observations by Ettori & Eckert (2022) for massive clusters.
@@ 4.5 Mass decomposition from the Euler equation
The approach in this section makes fewer assumptions than using the virial theorem. In virial equilibrium, the acceleration vanishes, and the remainder term $M_{\rm rem}$ would account only for those turbulent motions that are averaged out when computing the velocities in angular bins. Furthermore, for the computation of the hydrostatic mass, we assume spherical symmetry, hence the angular derivatives $\frac{\partial v_{r}}{\partial\theta}$ and $\frac{\partial v_{r}}{\partial\phi}$ vanish. For pure pressure support, the radial derivative of the velocity vanishes $\frac{\partial v_{r}}{\partial r}=0$, and there is no angular motion $v_{\theta}=v_{\phi}=0$. Taking all those narrowing assumptions together, the only remaining term is $M_{\rm therm}$, which is what is usually considered when computing hydrostatic equilibrium masses.
@@ 5 Conclusions
  • The hydrostatic mass bias is highly sensitive to the weighting applied to particles when computing the spherically averaged thermodynamic profiles (Fig. 2). Compared with X-ray luminosity weighted profiles, volume weighted profiles, which are theoretically preferable but cannot be obtained observationally, result in a much small hydrostatic mass bias for $M_{\text{500c}}\gg 10^{13}~{}\text{M}_{\odot}$. This suggests that much of the observed bias is not due to deviations from thermal hydrostatic equilibrium, but is caused by the biased nature of X-ray observations.
@@
By comparing different FLAMINGO simulations, we have shown that the hydrostatic mass bias is strongly dependent on the median cluster gas fractions and thus on the strength of stellar and AGN feedback. Because the differences between reasonable models is larger than the error bars on many cluster mass measurements, it would be advisable to include the model uncertainty when using observed cluster masses for cosmology. On top of that, the strong systematic dependence of the bias on the quality of the fit to the density and temperature profiles potentially creates another source of significant systematic uncertainty that should be accounted for.

### v2

@@ 1 Introduction
Alternatively, the mass can be estimated using the hydrostatic mass inferred from X-ray observations. Using the surface brightness distribution, the gas density profile can be extracted (e.g. Pratt & Arnaud, 2002; Ettori et al., 2013; Ghirardini et al., 2019). Assuming an emission model, the spatially resolved X-ray spectrum can be converted into a temperature profile (e.g. Arnaud, 1996). Under the assumption of hydrostatic equilibrium, fits to these profiles with an analytic function can be used to compute the enclosed mass (see e.g. Sarazin, 1988).
In practice, hydrostatic mass measurements pose a significant challenge. To obtain the mass enclosed within a sphere, the density and temperature profiles have to be three-dimensional, which means observations have to be deprojected from the plane of the sky in which they are observed (see e.g. Turner et al., 2024, for a comprehensive description of the process). Unlike in the case for weak lensing mass estimates, the cluster shape only has a 1 per cent effect on the mass estimate (Buote & Humphrey, 2012). However, temperature inhomogeneities within a single radial bin of the profile can have a 10 per cent effect on the mass, due to clumpy cold gas biasing the mean temperature low (Mazzotta et al., 2004; Rasia et al., 2014; Henson et al., 2017). Forward modelling and external observations, e.g. of the Sunyaev-Zeldovich effect, can improve the hydrostatic mass measurements and the estimates of the associated errors.
@@
The hydrostatic masses inferred from X-ray observations are thus found to be biased. One method of determining the hydrostatic mass bias is by comparing the X-ray inferred masses ($M_{\rm 500c,X}$) with weak lensing masses ($M_{\rm 500c,WL}$). Using this approach, the XXL cluster survey found a value of $M_{\rm X}/M_{\rm WL}=0.72\pm 0.08$ for $2\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ (Eckert et al., 2016), and the Canadian Cluster Project in combination with the Multi Epoch Nearby Cluster Survey obtained $M_{\rm X}/M_{\rm WL}=0.84\pm 0.04$ for $4\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<2\times 10^{15}~{}\mathrm{M_{\odot}}$ (Herbonnet et al., 2020), a slightly smaller bias than $M_{\rm X}/M_{\rm WL}=0.74\pm 0.06$ measured by Lovisari et al. (2020) who expanded the measurements with new X-ray and weak lensing observations in the mass range $3\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<2\times 10^{15}~{}\mathrm{M_{\odot}}$. However, much weaker biases of $M_{\rm X}/M_{\rm WL}=0.95\pm 0.05$ for $2\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ were found by Smith et al. (2016) and later $M_{\rm X}/M_{\rm WL}=0.84\pm 0.095$ for $3\times 10^{14}~{}\mathrm{M_{\odot}}<M_{\rm 500c}<10^{15}~{}\mathrm{M_{\odot}}$ by Wicker et al. (2023).
@@ 2.1 The FLAMINGO simulations
FLAMINGO (Full-hydro Large-scale structure simulations with All-sky Mapping for the Interpretation of Next Generation Observations) is a large suite of hydrodynamical cosmological simulations, covering large cosmic volumes. The flagship run comprises a region of $\mathrm{(2.8~{}Gpc)^{3}}$, which is sufficiently large for statistical studies of galaxy groups and clusters. At $z=0$ FLAMINGO contains $>10^{6}$ haloes with $M_{\rm 500c}\geq 10^{13}~{}\mathrm{M_{\odot}}$, $>10^{5}$ haloes with $M_{\rm 500c}\geq 10^{14}~{}\mathrm{M_{\odot}}$, and $461$ haloes with $M_{\rm 500c}\geq 10^{15}~{}\mathrm{M_{\odot}}$. The simulations are described in detail in Schaye et al. (2023). A unique feature is the machine learning-aided calibration of the subgrid prescriptions for stellar and AGN feedback to the observed $z=0$ galaxy stellar mass function and $z=0.1-0.3$ cluster gas fractions (Kugel et al., 2023). Variations on the fiducial model are made by shifting the observed gas fractions up and down by a multiple of their uncertainty ($\sigma$), and recalibrating the model to fit those new data points. FLAMINGO also includes models that vary the galaxy mass function or the cosmology. We will not use the latter here as we found that the cosmology variations do not affect the bias.
@@ 2.4 Fitting the thermodynamic profiles
where $\alpha$, $\beta$, $\gamma$, $\epsilon$ are free parameters, as well as the normalisation of the density profile $\rho(0)$, the core radius $r_{\rm c}$, and the scale radius $r_{\rm s}$. The scale radius ($r_{\rm s}$) allows for a second characteristic transition radius in the profile. Though this model has 7 free parameters, $\gamma$ is conventionally set to $\gamma=3$, which we adopt throughout this work. Note that because we only fit the region 0.5-2.0 $r_{\rm 500c}$, the parameters related to the core remain largely unconstrained.
@@
an eight-parameter function, where $T_{\rm 0}$ represents the temperature normalisation, $T_{\rm min}$ the minimum temperature of the core, $r_{\rm c}$ the core radius, $r_{\rm t}$ the transition radius, which, along with $\alpha$, $\alpha_{\rm c}$, $\beta$, and $\gamma$, are free parameters. Similar to the modified-$\beta$-profile, out choice to fit between 0.5-2.0 $r_{\rm 500c}$ leaves the parameters related to the core largely unconstrained.
@@ 3.1 Dependence on mass
Observational determinations of the bias, shown as the red data points, scatter widely below $10^{15}~{}\mathrm{M_{\odot}}$. We extract the data points from the observational literature by taking the highest and lowest mass in each observed sample and applying the power-laws, with errors, fitted by the respective authors at those masses. These data points are obtained using either $\beta$- or double-$\beta$-profiles to fit the density profile, and the Vikhlinin (2006) prescription for the temperature profile. Our results agree very well with Wicker et al. (2023) at $3\times 10^{14}~{}\mathrm{M_{\odot}}$, and have a very similar mass trend. In the study from Eckert et al. (2016) the inverse mass trend is found, where the bias decreases with increasing mass, though the magnitude is very similar to our result and the trend is weak. Herbonnet et al. (2020) find no clear mass trend with a slight dip in the bias around $10^{14.8}~{}\mathrm{M_{\odot}}$. We conclude that the FLAMINGO hydrostatic biases are in good agreement with the observations, but that the observational constraints are rather uncertain for all but the highest masses.
@@ 3.3 X-ray versus volume weighted averaging
Because hydrostatic masses are inferred from spherically averaged profiles, any deviations from spherical symmetry will complicate their interpretation. The presence of structure will lead to deviations between thermodynamic profiles computed using different weighting schemes. Our approach of using X-ray weighted thermodynamic profiles in the hydrostatic mass inference is appropriate when comparing with results from X-ray observations, but to measure the true dynamical state of the cluster volume weighting is the natural choice. If the gas is clumpy, then X-ray weighting will favour denser gas, provided it is not sufficiently dense to cool to low temperatures, because the emissivity scales with the square of the density. In Appendix A we study the impact of subhaloes on the hydrostatic bias, and find it to be relatively small.
We compare the hydrostatic mass bias measured from X-ray- and volume weighted profiles. In earlier work we compared with mass-weighted profiles, and found them to fall in-between the X-ray- and volume-weighted cases (Braspenning et al., 2023). From Fig. 2, it is clear that the X-ray-weighted profiles result in a significantly stronger, and more mass dependent bias. As a large mass bias is driven by shallow gradients, this indicates that the thermal pressure gradient in X-ray-weighted gas is shallower. We conclude that for $M_{\text{500c}}\gg 10^{13}\,\text{M}_{\odot}$ most of the hydrostatic mass bias is due to X-rays being a biased tracer of the gas rather than due to true deviations from hydrostatic equilibrium.
@@ 3.4 Model variations
We note that between $7\times 10^{14}~{}\mathrm{M_{\odot}}$ and $10^{15}~{}\mathrm{M_{\odot}}$, the bias inferred from observations shows significant scatter. It ranges from a bias exceeding that found for the highest gas fraction variation in FLAMINGO (Herbonnet et al., 2020, at $7\times 10^{14}~{}\mathrm{M_{\odot}}$) to a smaller bias than predicted by the lowest gas fraction variation (Wicker et al., 2023; Herbonnet et al., 2020, at $10^{15}~{}\mathrm{M_{\odot}}$). In general, the difference between the gas fraction variations is of the same order as the error on the observationally determined bias.
@@ 3.5 Redshift evolution
Between $z=0$ and $z=1.5$ the number of haloes decreases from $>$$10^{6}$ to $>$$10^{5}$ for $M_{\rm 500c}\geq 10^{13}~{}\mathrm{M_{\odot}}$, from $>$$10^{5}$ to $>$$10^{3}$ for $M_{\rm 500c}\geq 10^{14}~{}\mathrm{M_{\odot}}$, and from $461$ to 00 for $M_{\rm 500c}\geq 10^{15}~{}\mathrm{M_{\odot}}$.
@@ 4.1 Fit quality
Figure 7: The logarithmic derivatives of the luminosity-weighted modified-$\beta$ density profiles and the luminosity-weighted temperature profiles, at $r_{\rm HSE,500c}$, for the objects with the 5 per cent lowest hydrostatic masses (black) or 5 per cent highest masses (red) in 5 narrow true mass bins. The legend in each panel indicates the median bias values of each subset. The two sets of contours are consistently offset and can be connected by moving almost orthogonal to the thin black diagonal lines which correspond to a constant pressure derivative, showing that the difference in bias could be explained by differences in the profile derivatives at a constant temperature (see Eq. 2).
@@ 4.2 Drivers of outlying hydrostatic mass
Fig. 7 reveals a separation between the two groups in the $\mathrm{dlog}\rho$-$\mathrm{dlog}T$ plane, with the derivatives computed at $r_{\rm HSE,500c}$. Moving along the thin diagonal lines, the bias would change due to a different temperature at a fixed sum of derivatives (i.e. a fixed pressure derivative, see Eq. 2), whereas moving perpendicular to those lines the bias changes at fixed temperature, but due to changing derivatives.
@@
From Fig. 7 it can be seen that the two subsets can be connected by moving almost perpendicularly to the thin lines, indicating that the change in inferred hydrostatic mass could be explained by a change in the pressure gradient at fixed temperature. This agrees with results from Ansarifard et al. (2020), who provide a partial empircal correction to the hydrostatic mass bias by using the gradient of the density profile. To confirm this, we examine the temperatures at $r_{\rm HSE,500c}$ for these two groups of outliers. Fig. 8 illustrates that objects exhibiting low and high inferred hydrostatic masses have comparable temperatures within identical narrow true mass bins. Consequently, we infer that the scatter in the bias at a fixed mass and at fixed goodness-of-fit can be accounted for by differences in the derivatives of the temperature and density profiles, as depicted in Fig. 7, indicating steeper (shallower) pressure profiles for the largest (smallest) hydrostatic masses at fixed true mass.
@@ 4.4 Energy balance
Martizzi & Agrusa (2016), who simulated clusters in the $M_{\rm 500c}\approx 1-3\times 10^{14}~{}\mathrm{M_{\odot}}$ range, found that, using volume weighting, at $r_{\rm 500c}$ non-thermal energy can fully account for the hydrostatic bias. We find that in this mass range, the non-thermal energy in FLAMINGO is indeed sufficient to achieve virial equilibrium if we ignore external pressure. Our estimate of $\approx 10$ per cent non-thermal energy is furthermore in line with what is found in observations by Ettori & Eckert (2022) for massive clusters.
@@ 4.5 Mass decomposition from the Euler equation
The approach in this section makes fewer assumptions than using the virial theorem. In virial equilibrium, the acceleration vanishes, and the remainder term $M_{\rm rem}$ would account only for those turbulent motions that are averaged out when computing the velocities in angular bins. Furthermore, for the computation of the hydrostatic mass, we assume spherical symmetry, hence the angular derivatives $\frac{\partial v_{r}}{\partial\theta}$ and $\frac{\partial v_{r}}{\partial\phi}$ vanish. For pure pressure support, the radial derivative of the velocity vanishes $\frac{\partial v_{r}}{\partial r}=0$, and there is no angular motion $v_{\theta}=v_{\phi}=0$. Taking all those narrowing assumptions together, $M_{\rm rot}=M_{\rm stream}=M_{\rm accel}=0$, and the only remaining term is $M_{\rm therm}$, which is what is usually considered when computing hydrostatic equilibrium masses.
@@ 5 Conclusions
  • The hydrostatic mass bias is highly sensitive to the weighting applied to particles when computing the spherically averaged thermodynamic profiles (Fig. 2). Compared with X-ray luminosity weighted profiles, volume weighted profiles, which are theoretically preferable but cannot be obtained observationally, result in a much small hydrostatic mass bias for $M_{\text{500c}}\gg 10^{13}~{}\text{M}_{\odot}$. This suggests that much of the observed bias is not due to deviations from thermal hydrostatic equilibrium, but is caused by the biased nature of X-ray emitting gas.
@@
By comparing different FLAMINGO simulations, we have shown that the hydrostatic mass bias is strongly dependent on the median cluster gas fractions and thus on the strength of stellar and AGN feedback. Because the differences between reasonable gas fraction models is larger than the error bars on many cluster mass measurements, tests of observational methods should thus use hydrodynamical simulations that span a range of gas fractions. This is also critical for estimating errors on X-ray inferred masses for cluster cosmology. On top of that, the strong systematic dependence of the bias on the quality of the fit to the density and temperature profiles potentially creates another source of significant systematic uncertainty that should be accounted for.